if a family with only two daughters, one of the daughters is currently three times as old aa the other. two years ago the older daughter was four times as old as the younger sister was then. What is the current age of both daughters ?
Answer by ikleyn(52787)

.

Let x be the current age of the younger daughter.

Then the age of the oldest is 3x, according to the condition.


Two years ago the younger was  x-2 years old, while the older was 3x-2 years old.


The problem says


    4*(x-2) = 3x - 2.


Simplify step by step and find x


    4x - 8 = 3x - 2

    4x - 3x = 8 - 2

     x      = 6.


ANSWER.  The younger daughter is 6 years old now.  The older is 3*6 = 18 years old.


CHECK.  2 years ago they were 4 years and 16 years old, with the ratio of 4.    ! Correct !

Solved.

---------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.